### v1

@@ Abstract
[Abstract] Cloud-native applications are increasingly becoming popular in modern software design. Employing a microservice-based architecture into these applications is a prevalent strategy that enhances system availability and flexibility. However, cloud-native applications also introduce new challenges, such as frequent inter-service communication and the complexity of managing heterogeneous codebases and hardware, resulting in unpredictable complexity and dynamism. Furthermore, as applications scale, only limited research teams or enterprises possess the resources for large-scale deployment and testing, which impedes progress in the cloud-native domain. To address these challenges, we propose CloudNativeSim, a simulator for cloud-native applications with a microservice-based architecture. CloudNativeSim offers several key benefits: (i) comprehensive and dynamic modeling for cloud-native applications, (ii) an extended simulation framework with new policy interfaces for scheduling cloud-native applications, and (iii) support for customized application scenarios and user feedback based on Quality of Service (QoS) metrics. CloudNativeSim can be easily deployed on standard computers to manage a high volume of requests and services. Its performance was validated through a case study, demonstrating higher than 94.5% accuracy in terms of response time. The study further highlights the feasibility of CloudNativeSim by illustrating the effects of various scaling policies.
@@ 1 Introduction
In modern software development, the trend of cloud-native applications is on the rise. More and more applications are migrated to cloud-native environments, offering reliable and flexible functions. When choosing the development architecture for cloud-native products, a microservice-based architecture is considered more efficient compared to traditional monolithic architectures 1, 2. Figure 1 presents a demonstration of a cloud-native application based on microservice-based architecture, detailing the interactions between clients, microservices, and data centers. Clients send requests through application interfaces, which generate various service dependencies that are mapped to different instances (e.g., pods and containers). These microservice instances are deployed in suitable resource environments, supported by the resources of the data center, ensuring efficient and reliable service delivery. This architecture provides numerous advantages, including high availability and decentralization. These advantages can accelerate development and deployment processes, offer flexible auto-scaling mechanisms, and reduce maintenance costs 3, 4.
However, the transition to a microservice-based architecture faces challenges of complexity and dynamism 5, 6, 7. microservice-based architecture typically operates on heterogeneous code and hardware. An application built on a microservice-based architecture may consist of intricate modules and complex dependencies. Moreover, to offer adaptable virtualization for these modules, microservice-based architecture introduces numerous entities, such as requests and services. As the application scales, the communications between these entities become more frequent, requiring the coordination of scheduling policies, and leading to an uncontrollable dynamism.
Simultaneously, the high barriers to deploying large-scale cloud-native applications are also challenging for researchers. Large-scale applications require a robust and extensive data center capable of handling numerous requests and allocating sufficient resources to internal entities. Fewer teams have the capability to fulfill these requirements, let alone to assess scheduling policies based on microservice-based architecture.
@@
  • Comprehensive modeling approach: CloudNativeSim comprehensively models cloud-native applications and microservice-based architecture, clearly illustrating the dependency graph and resource transitions between entities.
  • High extensibility and customization: To ensure extensibility, our simulator inherits the core design of CloudSim but extends it into a new framework. Additionally, through the file registry function, users can easily customize different scenarios for their applications.
@@
  • Simulation of dynamic request generation and dispatching: CloudNativeSim features a built-in request generator to simulate the dynamic process of clients randomly generating requests and distributing them to different services.
  • Innovative cloudlet scheduling mechanism: During the simulator’s runtime, cloudlets will flow along the service dependency graph, undergoing various phases such as waiting, execution, and spawning under the supervision of a scheduler.
  • Latency calculation based on the critical path: The service dependency graph topology is characterized as a typical directed acyclic graph (DAG), which contains numerous parallel chains. CloudNativeSim will compute request latency by identifying the critical path within these chains.
  • QoS metrics feedback: Considering the quality of experience, the simulator also offers various optional feedback on metrics related to requests, displaying this information in the system logs.
  • New policy interfaces: Microservice-based architecture introduces a series of new policies, such as service allocation and scaling policies. CloudNativeSim implements some universal interfaces for these new policies.
@@
The key contributions of CloudNativeSim include: (i) a comprehensive modeling approach for cloud-native applications and microservice-based architecture, (ii) the ability to run on a standard personal computer while handling a large volume of requests and services concurrently, (iii) achieving a response time simulation accuracy of over 94.5%, and (iv) providing a range of scheduling policy interfaces, enabling customized implementation and evaluation of various scheduling policies.
@@
The structure of this paper is organized as follows: Section 2 presents a discussion of contemporary distributed system simulations and modeling, highlighting the unique capabilities of CloudNativeSim. Section 3 details CloudNativeSim’s simulation framework, emphasizing the functions of each architectural component. Section 4 describes the key components related to requests and cloudlets, achieving a comprehensive application modeling. Section 5 discusses the design of service scheduling policies, including resource allocation strategies and service scaling mechanisms. Finally, Section 6 provides comprehensive evaluations of CloudNativeSim, validating its simulation capacity and response time accuracy, and demonstrating the request generator and scaling algorithms. Section 7 concludes the paper with a brief summary and suggestions for future research directions.
@@ 2 Related Work
Simulator Names   Execution Environment   Datacenter Configuration   Resource Characteristic   Instance Management   ServiceChain Modeling   Service Allocation   Horizontal Scaling   Vertical Scaling   Service Migration   Tasks Scheduling   Request Processing   QoS Feedback
CloudSim 3.0 8
@@
CloudSim Plus 9 Cloud $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
NetworkCloudSim 10 Cloud $\checkmark$ $\checkmark$ $\checkmark$
iFogSim 11, 12 Fog/Edge $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ContainerCloudSim 13 Cloud $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ServiceSim 14 Cloud/Edge $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ColocationSim 15 Cloud $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PerfSim 16 Cloud/Edge $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
As cloud computing technology evolves, a growing array of simulators have been developed to support performance evaluations from cloud to edge environments. CloudSim 8 and its derivatives 9, 10, 11, 12, 13, 14 stand out as the principal tools in this domain. Based on a discrete-event simulation framework, CloudSim supports extensive data centers and energy-efficient computational resources. It introduces simulation events and entities, facilitating detailed modeling and analysis. CloudSim employs the concept of "cloudlets" to represent tasks, with their length indicating the number of instructions to be processed. These cloudlets can be managed by time-sharing or space-sharing methods to optimize the utilization of processing elements (PE) in virtual machines. After version 3.0, CloudSim has incorporated new features, including support for Software-Defined Networking (SDN) and containers. However, these enhancements have also increased the simulator’s complexity. Notably, CloudSim lacks inherent support for microservices, limiting its effectiveness in modeling modern cloud-native applications.
@@
CloudSim Plus 9, an iteration of CloudSim 3.0, is redesigned using Java 17 to reduce code duplication and promote code reuse, enhancing both scalability and precision. This iteration offers a cleaner, more functional, and user-friendly interface, and it introduces more advanced functions such as auto-scaling for virtual machines, heuristics for resource management, and lifetime control of entities. However, despite these advancements, like CloudSim, it does not provide support for microservice architectures.
@@
Following CloudSim’s foundational framework, numerous simulators have been developed. Among them, NetworkCloudSim 10, an extension of CloudSim, is specifically designed to simulate network communications in cloud computing environments. It incorporates a network model into CloudSim, allowing researchers and developers to account for factors such as bandwidth limitations and network latencies within and between data centers. This enhancement enables more precise simulation and evaluation of cloud computing applications’ performance. Notably, NetworkCloudSim introduces a TaskStage field in the cloudlet and considers the derivative relationships between cloudlets to a certain extent. Although NetworkCloudSim has enhanced network simulation, it lacks the capability to simulate communication between microservices, which is essential for modern cloud-native applications.
@@
iFogSim 11, a widely used extension of CloudSim, introduces a new framework that revolves around a controller. It focuses on the mobility of Internet of Things (IoT) devices and clustering mechanisms, allowing more accurate simulations of dynamic and distributed edge environments. In its second version 12, iFogSim extends to include a microservice-based architecture, incorporating features such as service discovery, placement, horizontal scaling, and load balancing. It also addresses the deployment and routing of microservices, which is crucial for system flexibility 17. However, iFogSim is primarily designed for Fog/Edge environments rather than cloud-native environments and has limitations in supporting microservice architectures, such as modeling service chains and resource scaling.
ContainerCloudSim 13, by introducing container technology, represents a significant innovation in CloudSim. It builds a detailed container abstraction on top of virtual machines and extends new strategies for container placement, selection, provisioning, and overbooking. Although ContainerCloudSim does not emphasize microservice-based architectures, its in-depth portrayal of containers provides a robust framework for modeling instances of services. This allows researchers to explore container-specific scenarios and optimizations, such as resource overbooking and efficient container provisioning, which are critical for modern cloud environments. However, the lack of support for microservices remains a limitation in fully capturing the dynamics of cloud-native applications.
ServiceSim 14, a simulator introduced recently, integrates the strengths of iFogSim and NetworkCloudSim. It not only facilitates service discovery and load balancing but also considers the relationship between user requests and cloudlets. Additionally, it incorporates NetworkCloudSim’s task stage mechanism to differentiate various stages of a request. ServiceSim places heightened focus on the process of client requests and service discovery. However, it lacks a detailed depiction of microservice-based architecture and falls short in offering new policy interfaces. The QoS feedback provided to users is also insufficient.
ColocationSim 15, a novel simulator, focuses on the collaborative deployment of latency-critical (LC) jobs and best-effort (BE) jobs. Instead of using a discrete event framework, it employs queuing theory models to calculate the response time and resource overhead of a microservice system. However, while queuing theory models are effective in many scenarios, their reliability and applicability in microservice architectures are not guaranteed. Microservices often exhibit complex interactions and dependencies that may not be adequately captured by traditional queuing models. Therefore, the practical deployment of ColocationSim in microservice-based environments requires careful consideration of these limitations, alongside ongoing validation against real-world data to ensure model accuracy and relevance.
PerfSim 16 is another simulator that differs from CloudSim. PerfSim focuses on accurately simulating complex dependencies between services, effectively capturing and analyzing the impact of these dependencies on system performance. It not only emphasizes the mathematical modeling of service dependencies but also supports both horizontal and vertical scaling to meet users’ performance optimization needs. Additionally, PerfSim emphasizes the processing of user requests and models the transmission of packets in the network. However, due to the lack of a discrete event mechanism, its flexibility is somewhat limited. This limitation means that during simulation, PerfSim may not respond as flexibly to different simulation scenarios and dynamic changes as other simulators that are equipped with a discrete event mechanism.
Compared with the aforementioned simulators, CloudNativeSim proposed in this paper has the following notable features. (i) It focuses on dynamic cloud-native environments while inheriting the basic modules of CloudSim. (ii) It clearly models service chains and entities which are the unique features of cloud-native applications. (iii) It provides extensible scheduling policy interfaces, such as service allocation and scaling to model resource scheduling process for cloud-native applications. (iv) It includes a dynamic task scheduling module for cloud-native applications. (v) It possesses a complete request handling process and QoS metric feedback. These features make CloudNativeSim both practical and extensible.

### v2

@@ Abstract
[Abstract] Cloud-native applications are increasingly becoming popular in modern software design. Employing a microservice-based architecture into these applications is a prevalent strategy that enhances system availability and flexibility. However, cloud-native applications introduce new challenges, including frequent inter-service communication and the management of heterogeneous codebases and hardware, resulting in unpredictable complexity and dynamism. Furthermore, as applications scale, only limited research teams or enterprises possess the resources for large-scale deployment and testing, which impedes progress in the cloud-native domain. To address these challenges, we propose CloudNativeSim, a simulator for cloud-native applications with a microservice-based architecture. CloudNativeSim offers several key benefits: (i) comprehensive and dynamic modeling for cloud-native applications, (ii) an extended simulation framework with new policy interfaces for scheduling cloud-native applications, and (iii) support for customized application scenarios and user feedback based on Quality of Service (QoS) metrics. CloudNativeSim can be easily deployed on standard computers to manage a high volume of requests and services. Its performance was validated through a case study, demonstrating higher than 94.5% accuracy in terms of response time simulation. The study further highlights the feasibility of CloudNativeSim by illustrating the effects of various scaling policies.
@@ 1 Introduction
In modern software development, the trend of cloud-native applications is on the rise. More and more applications are migrated to cloud-native environments, offering reliable and flexible functions. When choosing the development architecture for cloud-native products, a microservice-based architecture is considered more efficient compared to traditional monolithic architectures 1, 2. Figure 1 presents a demonstration of a cloud-native application based on microservice-based architecture, detailing the interactions between clients, microservices, and data centers. Clients send requests through application interfaces, which generate various service dependencies that are mapped to different instances (e.g., pods and containers). These microservices instances are deployed in suitable resource environments, supported by the resources of the data center, ensuring efficient and reliable service delivery. This architecture provides numerous advantages, including high availability and decentralization. These advantages can accelerate development and deployment processes, offer flexible auto-scaling mechanisms, and reduce maintenance costs 3, 4, 5.
However, the transition to a microservice-based architecture faces challenges of complexity and dynamism 6, 7, 8. A microservice-based architecture typically operates on heterogeneous code and hardware. An application built on a microservice-based architecture may consist of intricate modules and complex dependencies. Moreover, to offer adaptable virtualization for these modules, microservice-based architecture introduces numerous entities, such as requests and services. As the application scales, the communications between these entities become more frequent, requiring the coordination of scheduling policies, and leading to an uncontrollable dynamism.
Furthermore, researchers face significant challenges in deploying large-scale cloud-native applications. Large-scale applications require a robust and extensive data center capable of handling numerous requests and allocating sufficient resources to internal entities. Few teams have the capability to fulfill these requirements, let alone to assess scheduling policies based on microservice-based architecture.
@@
  • Comprehensive modeling approach: CloudNativeSim is a comprehensive modeling framework designed specifically for cloud-native applications and microservice-based architectures. By incorporating essential components such as services, pods or containers, virtual machines (VMs), and datacenters, it offers more comprehensive object modeling capabilities compared to existing simulators.
  • File registration mechanism: To address the inherent complexity of microservice architectures, which involve numerous configuration units like services and pods, CloudNativeSim implements a file-based registration system. This mechanism enables batch registration of architectural components through configuration files, significantly reducing the manual configuration burden and potential for errors.
@@
  • Dynamic request generation and distribution: CloudNativeSim features a built-in request generator that accurately simulates real-world scenarios. This component handles the dynamic process of clients randomly generating requests and intelligently distributing them across different services, allowing users to test various load patterns and service interactions.
  • Innovative cloudlet scheduling mechanism: Recognizing the limitations of traditional scheduling approaches in microservice environments, CloudNativeSim introduces an enhanced cloudlet scheduling mechanism. During runtime, cloudlets traverse through the service dependency graph under scheduler supervision, managing complex state transitions that include waiting, execution, and request propagation phases. This mechanism is specifically designed to handle the dynamic nature of microservice architectures, where requests flow through multiple interconnected services.
  • Critical path-based latency calculation: The simulator models service dependencies as a Directed Acyclic Graph (DAG), enabling detailed performance analysis. By identifying critical paths within parallel execution chains, CloudNativeSim provides accurate latency calculations that reflect real-world service interactions and potential bottlenecks.
  • Dynamic QoS metrics record: CloudNativeSim tracks and records dynamic changes in both system structure and QoS metrics during runtime. The QoS metrics include request-related measurements such as latency, Requests Per Second (RPS) for different Application Programming Interfaces (APIs), and Service Level Objective (SLO) violation rates as well as resource utilization metrics (CPU and memory usage) at both instance and VM levels. Through multiple channels including system logs and Grafana  [^1] https://grafana.com/ integration, these dynamic metrics can be monitored and visualized in real-time, providing comprehensive insights into system behavior and performance evolution.
  • Extensible policy interfaces: To accommodate the diverse requirements of microservice architectures, CloudNativeSim provides a comprehensive set of policy interfaces. These interfaces support various operational aspects including service allocation, scaling decisions, and resource management, allowing users to implement and evaluate customized policies tailored to their specific needs.
@@
The key contributions of this work include: (i) a comprehensive modeling approach for cloud-native applications and microservice-based architecture, (ii) the ability to run on a standard personal computer while handling a large volume of requests and services concurrently, (iii) achieving a response time simulation accuracy of over 94.5%, and (iv) providing a range of scheduling policy interfaces, enabling customized implementation and evaluation of various scheduling policies.
@@
The structure of this paper is organized as follows: Section 2 presents a discussion of contemporary distributed system simulations and modeling, highlighting the unique capabilities of CloudNativeSim. Section 3 details CloudNativeSim’s simulation framework, emphasizing the functions of each architectural component. Section 4 describes the key components related to requests and cloudlets, presenting advanced application modeling. Section 5 discusses the design of service scheduling policies, including resource allocation strategies and service scaling mechanisms. Finally, Section 6 provides comprehensive evaluations of CloudNativeSim, validating its simulation capacity and response time accuracy, and demonstrating the request generator and scaling algorithms. Section 7 concludes the paper with a brief summary and suggestions for future research directions.
@@ 2 Related Work
Simulator Names   Execution Environment   Datacenter Configuration   Resource Characteristic   Instance Management   ServiceChain Modeling   Service Allocation   Horizontal Scaling   Vertical Scaling   Service Migration   Tasks Scheduling   Request Processing   QoS Metrics
CloudSim 3.0 9
@@
CloudSim Plus 10 Cloud $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
NetworkCloudSim 11 Cloud $\checkmark$ $\checkmark$ $\checkmark$
iFogSim 12, 13 Fog/Edge $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ContainerCloudSim 14 Cloud $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ServiceSim 15 Cloud/Edge $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ColocationSim 16 Cloud $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PerfSim 17 Cloud/Edge $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
As cloud computing technology evolves, a growing array of simulators have been developed to support performance evaluations from cloud to edge environments. CloudSim 9 and its derivatives 10, 11, 12, 13, 14, 15 stand out as the principal tools in this domain. Based on a discrete-event simulation framework, CloudSim supports extensive data centers and energy-efficient computational resources. It introduces simulation events and entities, facilitating detailed modeling and analysis. CloudSim employs the concept of "cloudlets" to represent tasks, with their length indicating the number of instructions to be processed. These cloudlets can be managed by time-sharing or space-sharing methods to optimize the utilization of processing elements (PE) in virtual machines. After version 3.0, CloudSim has incorporated new features, including support for Software-Defined Networking (SDN) and containers. However, the fundamental cloudlet-based task model poses significant challenges for microservice simulation. Specifically, the basic structure of cloudlets lacks the flexibility to represent the dynamic nature of microservice interactions, service dependencies, and fine-grained resource requirements characteristic of microservice architectures. Given CloudSim’s extensible architecture that supports the customization of new entities and events, CloudNativeSim and various other derivatives have been developed upon this foundation, each specializing in distinct aspects of cloud computing simulation.
@@
CloudSim Plus 10, an iteration of CloudSim 3.0, is redesigned using Java 17 to reduce code duplication and promote code reuse, enhancing both scalability and precision. This iteration offers a cleaner, more functional, and user-friendly interface, and introduces more advanced functions such as auto-scaling for virtual machines, heuristics for resource management, and lifetime control of entities. Despite these improvements, CloudSim Plus inherits the fundamental limitations of CloudSim’s cloudlet-based architecture. The auto-scaling mechanisms, while suitable for virtual machines, are not designed to handle the rapid scaling and complex dependencies of microservices.
@@
Following CloudSim’s foundational framework, numerous simulators have been developed. Among them, NetworkCloudSim 11, an extension of CloudSim, is specifically designed to simulate network communications in cloud computing environments. It incorporates a network model into CloudSim, allowing researchers and developers to account for factors such as bandwidth limitations and network latencies within and between data centers. This enhancement enables more precise simulation and evaluation of cloud computing applications’ performance. NetworkCloudSim introduces a TaskStage field in the cloudlet and considers the derivative relationships between cloudlets, providing a foundation for modeling task dependencies. While these features are valuable for traditional distributed applications, the current implementation focuses primarily on network-level communication rather than service-level interactions that characterize microservice architectures.
@@
iFogSim 12, a widely used extension of CloudSim, introduces a controller-centric framework that focuses on the mobility of Internet of Things (IoT) devices and clustering mechanisms. Its second version 13 extends to include microservice-based features, such as service discovery, placement, horizontal scaling, and load balancing. It also addresses the deployment and routing of microservices, which is crucial for system flexibility 18. However, being primarily designed for Fog or Edge environments, its architecture and resource management mechanisms are optimized for edge computing scenarios rather than cloud-native environments, where different constraints and requirements may apply.
ContainerCloudSim 14, by introducing container technology, represents a significant innovation in CloudSim. It builds a detailed container abstraction on top of virtual machines and extends new strategies for container placement, selection, provisioning, and overbooking. Although ContainerCloudSim does not emphasize microservice-based architectures, its in-depth portrayal of containers provides a robust framework for modeling instances of services. However, extending ContainerCloudSim to support microservice architectures presents considerable challenges. First, its extensive code overrides of CloudSim’s base implementation create architectural constraints. More importantly, its fundamental design follows a "VM-to-containers" one-to-many mapping paradigm, which makes it difficult to implement the many-to-many "services-to-instances" relationships characteristic of microservice architectures. Adapting ContainerCloudSim to support such relationships would require substantial modifications to its core architecture and significantly increase code complexity.
ServiceSim 15, while representing an advancement in cloud simulation by integrating features from iFogSim and NetworkCloudSim, demonstrates the common limitations faced by current simulators. ServiceSim successfully implements essential functionalities such as service discovery and load balancing, while incorporating sophisticated task staging mechanisms. However, it still leaves gaps in comprehensive microservice modeling and policy interface flexibility. These limitations reflect the broader challenge of adapting existing simulation frameworks to fully support modern microservice architectures.
ColocationSim 16 exemplifies an alternative approach to microservice simulation through queuing theory models. While innovative in its handling of latency-critical and best-effort jobs, this mathematical approach highlights the inherent difficulties in accurately modeling microservice behaviors. The simulator reliance on traditional queuing models, though computationally efficient, struggles to capture the dynamic nature and complex interdependencies characteristic of microservice architectures. This limitation underscores the broader challenge of finding an appropriate balance between model simplicity and accurate representation of microservice complexity.
PerfSim 17 offers valuable insights through its distinctive approach to cloud simulation. By incorporating discrete event simulation mechanisms, it excels in modeling and simulating complex service dependencies, with particular strength in analyzing their performance implications. The simulator demonstrates sophisticated capabilities in mathematical modeling of service dependencies and supports both horizontal and vertical scaling mechanisms for performance optimization. Its robust handling of user requests and network packet transmission simulation provides a solid foundation for performance analysis. However, PerfSim’s primary focus on service chain performance modeling, rather than comprehensive cloud-native applications, limits its broader applicability.
In summary, while various cloud computing simulators have emerged to address different aspects of cloud and edge computing environments, they exhibit distinct limitations in their ability to comprehensively simulate modern microservice architectures. Traditional simulators like iFogSim and NetworkCloudSim, while robust in their respective focused areas, present significant challenges when attempting substantial modifications to support microservice-specific features. While specialized microservice simulators such as ServiceSim and PerfSim exist, their architectural approaches differ fundamentally from what is needed to fully capture the nature of cloud-native applications. Current limitations in microservice architecture modeling include incomplete entity representation, simplified request arrival patterns, and basic service scheduling strategies.
To address these systemic limitations and advance the field of cloud-native simulation, we propose CloudNativeSim with the following design objectives: (i) File Registration Mechanism: To manage the increased complexity introduced by microservice architectures, which involve numerous units such as services and Pods, CloudNativeSim implements a file-based batch registration system that significantly simplifies configuration management. (ii) Dynamic Request Generation and Distribution: The simulator features a built-in request generator that accurately models real-world scenarios by simulating how clients dynamically generate and distribute requests across various services. (iii) Enhanced Cloudlet Scheduling: Recognizing the limitations of traditional cloudlet scheduling in handling microservice architectures, CloudNativeSim introduces an advanced scheduling mechanism specifically designed to address the dynamic and complex nature of these systems. (iv) Delay Analysis and Critical Path Detection: CloudNativeSim models service dependencies as Directed Acyclic Graphs (DAGs), enabling precise calculation of request delays and identification of critical paths within parallel service chains. (v) Comprehensive QoS Metrics: The simulator provides extensive quality-of-service monitoring capabilities, offering detailed request-related feedback through system logs and integration with Grafana for enhanced visualization and analysis. and (vi) Lightweight and Extensible Policy Framework: To accommodate the diverse requirements of microservice architectures, CloudNativeSim implements flexible and lightweight interfaces for various policies, including service allocation and scaling, ensuring adaptability to different deployment scenarios.